OpenStudy (anonymous):

Medal&Fan Which ratio is NOT equal to 4/10? Select one: a. 2/5 b. 8/20 c. 12/18 d. 24/60

OpenStudy (anonymous):

c. 2/3 in the reduced form.

OpenStudy (anonymous):

alright. cool.(;

OpenStudy (anonymous):

\[\text{Simplify}\left[\left\{\frac{2}{5},\frac{8}{20},\frac{12}{18},\frac{24}{60}\right\}\right]=\left\{\frac{2}{5},\frac{2}{5},\frac{2}{3},\frac{2}{5}\right\} \]Thank you for the medal.